Let ω0<π/2. The value of ∑[sin(ω0n)/πn]^4 equals
1
Expert's answer
2019-09-18T12:04:55-0400




"\displaystyle\sum_{i=1}^ \infin{(sin(nx))^4 \over n^4}\approx1.015(sinx)^{69/20}, 0<x<\pi/2"


"\displaystyle\sum_{i=1}^ \infin{(sin(n\omega_0))^4 \over n^4}\approx1.015(sin(\omega_0))^{69/20}, 0<\omega_0<\pi/2"

Then


"\displaystyle\sum_{i=1}^ \infin{(sin(\omega_0n))^4 \over (\pi n)^4}\approx{1.015 \over \pi^4}(sin(\omega_0))^{69/20}, 0<\omega_0<\pi/2"